OpenStudy (anonymous):

1. Completely factor 4x2 + 12x + 8. The answer would be turning out to look something like this 4(x + ____)(x + ____)

OpenStudy (anonymous):

it would be awesome if you guys could help ! c:

OpenStudy (anonymous):

Okay lets do it step by step. We have\[4x^{2} +12x + 8\]Can you see that 4 is a factor of all of the terms? if I factor the 4, I get\[4(...)\]What goes in the brackets?

OpenStudy (anonymous):

would it be x^2?

OpenStudy (anonymous):

That is correct for the first term, but I am factoring from all of the terms So based on what you said we now have\[4x^{2} + 12x+8 = 4(x^{2}+...)\] What's missing?

OpenStudy (anonymous):

Oh it would be the 12x and the 8? but it would be as 12 without the x and a regular 8 as it was?

OpenStudy (anonymous):

Not quite, when I say I am factoring out 4 from each term, what I mean is I am dividing each term by 4, then putting the four outside of the bracket, and putting what's left after I divide inside the bracket Another way of thinking about it is "4 times what gives me: \[4x^{2}+12x+8"\] ie \[4(...)=4x^{2}+12x+8\]

OpenStudy (anonymous):

oh so 4 times 3 and 4 times 2 ?

OpenStudy (anonymous):

what do you do after you figure out that ?

OpenStudy (anonymous):

Okay so you understand why we have\[4(x^{2}+3x+2)\]So far?

OpenStudy (anonymous):

right !

OpenStudy (anonymous):

Do you have any idea how to go about factoring\[x^{2}+3x+2?\]

OpenStudy (anonymous):

very vaguely but I think I can go from there hopefully

OpenStudy (anonymous):

Give it a try and let me know if you need any help

OpenStudy (anonymous):

Ok awesome thanks so much !! :D

OpenStudy (anonymous):

Just for a check,it would turn out to be x(x+2)(x+1) right?

OpenStudy (texaschic101):

4(x^2 + 3x + 2) = 4(x + 2)(x + 1)

OpenStudy (anonymous):

Ok awesome thanks !